OpenStudy (girlstudy):

Change the speed 0.200 cm/s to units of kilometers per year.

OpenStudy (girlstudy):

Can Someone explain this step by step to me? I wasnt really well explained on how to this properly

OpenStudy (yanasidlinskiy):

Would you like to see it worked out?

OpenStudy (girlstudy):

Yes, that would be appreciated

OpenStudy (yanasidlinskiy):

Does that make sense? @agent0smith

OpenStudy (agent0smith):

Looks about right, just hard to read

OpenStudy (yanasidlinskiy):

EPIC FAIL ok.. Hold on..

OpenStudy (yanasidlinskiy):

Any better? Best I could get it to lol

OpenStudy (agent0smith):

Sure lol... wouldn't it be easier to just use the equation editor here? :P

OpenStudy (yanasidlinskiy):

Ohhh truuueee hahha. So dumb lol

OpenStudy (girlstudy):

@YanaSidlinskiy can you explain it please? im a bit confused

OpenStudy (agent0smith):

\[\large \frac{ 0.2 cm }{ 1 s} *\frac{ 1m }{ 100cm }*\frac{ 1km }{1000m }*\frac{ 60s }{ 1 \min }*\frac{ 60\min }{ 1 hr }*\frac{ 24 hr }{ 1day }*\frac{ 365 days}{ 1 year }\]

OpenStudy (agent0smith):

It's just a lot of unit conversions... first cm to km, and then seconds to years.

OpenStudy (girlstudy):

@agent0smith THank youuu!!